OpenStudy (anonymous):

using law of sines, solve: A=75 degrees, B=50 degrees, a=7 solve for the other three.

OpenStudy (anonymous):

sin(75)/7 = sin(50)/b therefore b = 5.55 C = 180 - A - B = 55 degrees sinC/c = sin(75)/7 therefore c = 5.94